A pedestrian waiting for the light to change at an intersection hears a car approaching with its horn blaring. The car's horn produces sound with a frequency of 381 Hz, but the pedestrian hears a frequency of 386 Hz . How fast is the car moving?
Answer #1

Speed of sound is 340 m/s

From Doppler equation

f=f_{o}\left ( \frac{V}{V-V_{s}} \right )

386=381[340/(340-Vs)]

340-Vs=335.6

Vs=4.4 m/s